### v3

# No-Regret Learning in Stackelberg Games with an Application to Electric Ride-Hailing

Abstract
We consider the problem of efficiently learning to play single-leader multi-follower Stackelberg games when the leader lacks knowledge of the lower-level game. Such games arise in hierarchical decision-making problems involving self-interested agents. For example, in electric ride-hailing markets, a central authority aims to learn optimal charging prices to shape fleet distributions and charging patterns of ride-hailing companies. Existing works typically apply gradient-based methods to find the leader’s optimal strategy. Such methods are impractical as they require that the followers share private utility information with the leader. Instead, we treat the lower-level game as a black box, assuming only that the followers’ interactions approximate a Nash equilibrium while the leader observes the realized cost of the resulting approximation. Under kernel-based regularity assumptions on the leader’s cost function, we develop a no-regret algorithm that converges to an $\epsilon$-Stackelberg equilibrium in $\mathcal{O}(\sqrt{T})$ rounds. Finally, we validate our approach through a numerical case study on optimal pricing in electric ride-hailing markets.
I Introduction
Many real-world systems, including traffic networks, communication systems, and smart grids, involve multiple self-interested agents that interact repeatedly. Such independent and self-interested decision-making may result in inefficient and socially undesirable outcomes due to misaligned objectives and a lack of coordination among the agents [dubey1986]. Therefore, from a global perspective, it is important to design intervention mechanisms such as tolls, pricing schemes, and subsidies that can steer agents towards a socially desirable and efficient behavior.
A prominent example of inefficient decision-making arises in ride-hailing markets, where service providers may flock to regions of high demand in search of passengers, leaving other customer locations underserved [BEOJONE2021102890]. Interestingly, the growing integration of electric vehicles (EVs) into ride-hailing fleets of companies such as Uber, Lyft, and Curb, introduces a lever for external intervention via electricity pricing [maljkovic2023hierarchical, Li2023taxi, malkovic2023learninghow]. A central authority, such as the government or power providers, may set spatially varying prices to shape fleet distribution and charging patterns to match demand and supply across regions.
Introducing a central authority to design intervention mechanisms imposes a bi-level structure that naturally fits into the framework of a single-leader, multi-follower Stackelberg game [von1952theory]. In this setting, the incentive designer assumes the role of the leader, while the self-interested agents act as followers. In the context of ride-hailing, the government or the power provider would act as the leader, while the competing ride-hailing fleets would serve as the followers. At the upper level, the leader selects incentives to minimize a societal cost, anticipating the followers’ best responses. At the lower level, the followers are considered rational agents that play a non-cooperative game, the outcome of which forms a Nash equilibrium. This paper investigates whether the leader can efficiently learn to solve Stackelberg games solely from observations of the societal cost to tackle the incentive design problem described above.
Existing approaches usually rely on first-order methods [grontas24, maljkovic2024decentralizedcomputationleadersstrategy, liu2022inducing, ji2021bilevel, goyal2023algorithm] or problem-specific solutions [l-se]. First-order methods typically estimate the gradient of the leader’s objective, i.e., the so-called hypergradient, but this estimation requires the leader to have access to estimates of both the lower-level Nash equilibrium and the gradient of the followers’ utility functions [grontas24]. While these methods may be suitable for bilevel optimization problems [anandalingam1992hierarchical], they become impractical when the lower-level involves self-interested agents that are unwilling to share their private information. In particular, this is evident in electric ride-hailing markets, where agents behave strategically and are unlikely to disclose sensitive information, such as the gradient of their utility functions. As a result, the lower-level problem effectively becomes a black box to the leader, making traditional approaches difficult to apply.
Several past works [sessa2020learning, maheshwari2023follower, zhao2023online] allow the leader to probe the followers with different incentives and observe the realized societal cost. The work of [sessa2020learning] focuses on a single-leader, single-follower setup, where the leader observes the follower’s noisy best response. Assuming that the follower’s response function lies in a reproducing kernel Hilbert space (RKHS), they propose a no-regret algorithm that converges to an approximately optimal incentive in $\mathcal{O}(\sqrt{T})$. In contrast, [maheshwari2023follower] considers a different approach to tackle multi-follower Stackelberg games based on zeroth-order estimation of the leader’s hypergradient. While their algorithm converges to an approximately stationary point in $\mathcal{O}(\sqrt{T})$, to ensure that it is an approximate Stackelberg equilibrium, they require a hard-to-verify assumption on the Hessian of the leader’s cost function. Moreover, their method estimates the hypergradient by probing followers with an incentive and a perturbed version thereof, which may be impractical when a central authority cannot evaluate multiple incentives simultaneously.
In this paper, we consider a class of multi-follower Stackelberg games for which our contributions are threefold:
  • We propose a novel no-regret algorithm for the leader to learn in Stackelberg games that leverages Gaussian process regression, assuming that the leader’s cost function satisfies the RKHS assumption.
  • We show that with high probability our algorithm converges in $\mathcal{O}(\sqrt{T})$ to an $\epsilon$-Stackelberg equilibrium. Our method requires no prior knowledge of the lower-level game and only assumes that the followers can approximate a Nash equilibrium within a number of rounds polynomial in $T$.
  • We demonstrate the applicability of our setup to electric ride-hailing markets and validate our algorithm through a numerical case study in this domain.
Notation: Let $\mathds{R}$ and $\mathbb{Z}_{(+)}$ denote the sets of real and (non-negative) integer numbers. For any $T\in\mathbb{Z}_{+}$, we let $T]=\{1,2,...,T\}$. Let $\mathds{1}$ be the vector of all ones. If $\mathcal{A}$ is a finite index, we let $x=(x_{i})_{i\in\mathcal{A}}$ be the concatenation of vectors $x_{i}$. For a real-valued function $f:\mathds{R}^{d}\rightarrow\mathds{R}$, we let $\nabla_{x}f=(\frac{\partial f}{\partial x_{i}})_{i=1}^{d}$ be its gradient.
II Problem setup
We consider a Stackelberg game with $N+1$ agents consisting of a leader $L$ and a set of followers denoted by $\mathcal{N}$. In a Stackelberg game, the leader first chooses an action $\pi\in\Pi$ from its action set $\Pi\subseteq\mathds{R}^{p}$. The followers then simultaneously respond to the leader’s action $\pi$ by choosing an action $x_{i}\in\mathcal{X}_{i}\subseteq\mathds{R}^{d}$, where $\mathcal{X}_{i}$ denotes the action set of follower $i\in\mathcal{N}$. Furthermore, the leader has a cost function given by $J:\Pi\times\mathcal{X}\rightarrow\mathds{R}$ and each follower has a utility function given by $U_{i}:\mathcal{X}\times\Pi\rightarrow\mathds{R}$, where $\mathcal{X}=\Pi_{i=1}^{N}\mathcal{X}_{i}$ denotes the action space of the joint action profile $x\coloneqq(x_{i})_{i\in\mathcal{N}}$.
II-A Lower-level game
Given the leader’s action $\pi$, at the lower level, the followers decision-making problem can be cast as a game $\Gamma(\mathcal{N},\{\mathcal{X}_{i}\}_{i=1}^{N},\{U_{i}\}_{i=1}^{N};\pi)$. Each agent aims to maximize its utility given the actions of the other agents and the action of the leader. A popular solution concept in such games is the Nash equilibrium at which no agent has an incentive to unilaterally deviate from its action.
Definition 1
For a given $\pi$, the joint action profile $x^{*}\in\mathcal{X}$ is an $\epsilon$-Nash equilibrium if $U_{i}(x_{i},x_{-i}^{*};\pi)\leq U_{i}(x_{i}^{*},x_{-i}^{*};\pi)+\epsilon$ holds for all $i\in\mathcal{N}$, where $x_{-i}\coloneqq(x_{j})_{j\in\mathcal{N}\setminus\{i\}}$ denotes the joint action of all agents except $i$. The action profile $x^{*}$ is a Nash equilibrium if $\epsilon=0$.
In the following, we assume that the game $\Gamma$ is concave and strongly monotone. Concavity ensures that a Nash equilibrium exists [rosen1965existence, Theorem 1], while strong monotonicity ensures that it is also unique [rosen1965existence, Theorem 2].
Assumption 1
$\Gamma(\mathcal{N},\{\mathcal{X}_{i}\}_{i=1}^{N},\{U_{i}\}_{i=1}^{N};\pi)$ is a concave game for every $\pi\in\Pi$. Namely, for each agent $i\in\mathcal{N}$ the set $\mathcal{X}_{i}$ is non-empty, compact, and convex, and the utility function $U_{i}(x_{i},x_{-i};\pi)$ is continuously differentiable in $x$, and concave in $x_{i}$ for all $x_{-i}\in\Pi_{j\neq i}\mathcal{X}_{j}$ and all $\pi\in\Pi$.
Assumption 2
$\Gamma(\mathcal{N},\{\mathcal{X}_{i}\}_{i=1}^{N},\{U_{i}\}_{i=1}^{N};\pi)$ is $\alpha$-strongly monotone for every $\pi\in\Pi$. Namely, the game pseudogradient $v:\mathds{R}^{Nd}\times\Pi\rightarrow\mathds{R}^{Nd}$ defined as:
$v(x;\pi)=(v_{i}(x;\pi))_{i=1}^{N},\ \text{where}$
$v_{i}(x;\pi)=-\nabla_{x_{i}}U_{i}(x_{i},x_{-i};\pi),\quad\forall x\in\mathcal{X},\forall i\in\mathcal{N},$
satisfies $\langle v(x;\pi)-v(x^{\prime};\pi),x-x^{\prime}\rangle\geq\alpha\left\lVert x-x^{\prime}\right\rVert^{2}$ for all $x,x^{\prime}\in\mathcal{X}$ and $\alpha>0$.
II-B Stackelberg game
Under Assumptions 1 and 2, a Stackelberg game can be expressed as follows:
Attheupperlevel,theleaderaimstominimizeitscost
[…]
III Motivating example: Ride-hailing markets
We formulate charging management in ride-hailing markets as a Stackelberg game. At the lower level, each ride-hailing company $i\in\mathcal{N}$ operates a fleet of $M_{i}\in\mathbb{Z}$ electric vehicles (EVs) to serve demand across multiple districts, each equipped with its own charging infrastructure. For a given pricing vector $\pi\in\mathds{R}^{d}$, companies aim to maximize their profit by optimally dispatching and later recharging their vehicles [maljkovic2023hierarchical]. Specifically, each company $i$ decides how many EVs to allocate to each of the $d$ districts, choosing
$x_{i}\in[0,x_{i,\max}]^{d}\cap\{x_{i}\in\mathds{R}^{d}\ \lvert\ \mathds{1}^{\top}x_{i}\leq M_{i}\}.$
Given the price vector $\pi$, we assume company $i$’s utility function follows the market share acquisition as in [10591280, bravo2018bandit]:
$U_{i}(x_{i},x_{-i};\pi)=\sum_{m=1}^{d}W_{m}\frac{x_{i,m}}{\sum_{j\in\mathcal{N}}x_{j,m}+\Delta_{m}}-x_{i,m}\pi_{m}.$ (1b)
The utility reflects the difference between the company’s revenue from serving demand across $d$ districts and recharging costs, where $0<W_{\min}\leq W_{m}\leq W_{\max}$ represents the total revenue potential of district $m\in[d]$. In high-demand regions, a larger number of service vehicles is required to meet the increased volume of requests and to prevent customer abandonment due to long waiting times. This is captured by the exogenous parameter $0<\Delta_{\min}\leq\Delta_{m}\leq\Delta_{\max}$, which accounts for the fraction of total revenue potential not being distributed among the ride-hailing companies. For utilities given by (1b), it can be verified that the game’s pseudogradient is strongly monotone [bravo2018bandit, Appendix A.2.].
At the upper level, the central authority seeks to guide the overall EV allocation towards a predefined target distribution, with the goals of reducing congestion, ensuring equitable service coverage across the city, and balancing grid demand. For instance, charging prices in remote areas may be set lower than in the central districts, encouraging ride-hailing companies to charge their EVs there. Formally speaking, since the central authority lacks information about the absolute number of operating EVs, we assume it sets charging prices $\pi\in[0,\pi_{\max}]^{d}$ to incentivize ride-hailing companies to match a target distribution $\xi^{*}\in\{\xi\in[0,1]^{d}\mid\mathds{1}^{\top}\xi=1\}$. The interaction between the companies results in a joint allocation of EVs $x(\pi)=(x_{1}(\pi),\ldots,x_{N}(\pi))$, allowing the central authority’s cost to be expressed as:
$J(\pi)=\biggl\lVert\frac{\sum_{i\in\mathcal{N}}x_{i}(\pi)}{\mathds{1}^{\top}\sum_{i\in\mathcal{N}}x(\pi)}-\xi^{*}\biggr\lVert^{2},$ (1c)
which quantifies the deviation between the achieved distribution of EVs and the desired one.
In reality, the central authority cannot compute the unique Nash equilibrium $x(\pi)$ without knowledge of the companies’ utility functions. Instead, it can set charging prices and wait while the companies repeatedly compete with each other for demand until the market stabilizes at an approximation of the unique Nash equilibrium. The central authority can then assess how closely the actual distribution of EVs aligns with its target distribution and adjust the prices accordingly. In the following, we present an algorithm that lets the leader learn an $\epsilon$-Stackelberg equilibrium, which is also applicable to our motivating example.
IV Learning a Stackelberg equilibrium
We consider the setting where the Stackelberg game is repeated over several rounds. In each round $t$, the leader selects an action $\pi^{t}$ that is observed by the followers. Then, in a subroutine, the followers aim to learn an approximation $x^{t}(\pi^{t})$ of the unique Nash equilibrium $x^{*}(\pi^{t})$. In the following, we treat the lower-level game as a black-box and merely assume that after finitely many iterations within a subroutine, $x^{t}(\pi^{t})$ is indeed an approximation of $x^{*}(\pi^{t})$.
We now focus on the leader who seeks to learn a Stackelberg equilibrium in a repeated fashion. As $\pi^{t}$ changes across rounds, the leader observes a sequence of time-varying costs. This motivates us to define the following notion of regret:
$R^{T}=\sum_{t=1}^{T}J(\pi^{t},x^{t}(\pi^{t}))-\min_{\pi\in\Pi}J(\pi,x^{*}(\pi)).$ (1d)
It measures the leader’s additional cost beyond its optimal value due to 1) not knowing $\pi^{*}\in\arg\min_{\pi\in\Pi}J(\pi,x^{*}(\pi))$ beforehand and 2) the followers learning an approximation $x^{t}(\pi^{t})$ of the unique Nash equilibrium $x^{*}(\pi^{t})$ given $\pi^{t}$. The leader has no-regret if $R^{T}/T\rightarrow 0$ as $T\rightarrow\infty$. As we will show in Theorem 1, presented later in this section, having no-regret implies convergence to a Stackelberg equilibrium.
To ensure that the leader is able to attain no-regret, we make the following assumptions. At the end of round $t$, the leader receives feedback information that it can use to update its action $\pi^{t+1}$ and thus improve its cost. A realistic feedback model is bandit feedback [Srinivas2010], where the leader merely observes its realized cost. In ride-hailing applications, this is satisfied since the central authority can assess how closely the actual distribution of EVs matches the target distribution.
Assumption 3
In round $t$, the leader observes the realized cost $J(\pi^{t},x^{t}(\pi^{t}))=J(\pi^{t},x^{*}(\pi^{t}))-\epsilon^{t}$, where $\epsilon^{t}$ is the error due to the followers playing an approximation of the unique Nash equilibrium $x^{*}(\pi^{t})$.
To ensure that the error $\epsilon^{t}$ is bounded, we assume that the leader’s cost function is Lipschitz-continuous with respect to the followers’ joint action.
Assumption 4
The cost function $J:\Pi\times\mathcal{X}\rightarrow\mathds{R}$ is $L_{J}$-Lipschitz continuous in $x\in\mathcal{X}$, i.e., $\lvert J(\pi,x)-J(\pi,\overline{x})\rvert\leq L_{J}\lVert x-\overline{x}\rVert$ holds for all $x,\overline{x}\in\mathcal{X}$ and $\pi\in\Pi$.
Since $\mathcal{X}$ is compact, we can define $M=\max_{x,\overline{x}\in\mathcal{X}}\lVert x-\overline{x}\rVert$. Then, Assumption 4 implies that $\lvert J(\pi,x)-J(\pi,\overline{x})\rvert\leq L_{J}M$ for all $x,\overline{x}\in\mathcal{X}$ and $\pi\in\Pi$.
Attaining no-regret is impossible without any regularity assumptions on the cost function [Srinivas2010]. In this work, we further assume that similar inputs lead to similar outputs. This is satisfied, for example, in ride-hailing applications, where similar charging prices set by the central authority lead to similar vehicle distribution patterns, implying similar costs for the central authority.
Assumption 5
The cost function $J:\Pi\rightarrow\mathds{R}$ has a compact domain $\Pi$. [^1] Note that at the Nash equilibrium $x^{*}(\pi)$ the leader’s cost $J$ is uniquely defined by $\pi$. Thus, with a slight abuse of notation, we can adopt a modified mapping $J:\Pi\rightarrow\mathds{R}$ with $J(\pi)=J(\pi,x^{*}(\pi))$. Furthermore, $J$ has a bounded norm $\|J\|_{k}=\sqrt{\langle J,J\rangle_{k}}\leq B$ in a reproducing kernel Hilbert space (RKHS, see [rasmussen2005]) associated with a positive semi-definite kernel function $k(\cdot,\cdot)$. The RKHS is denoted by $\mathcal{H}_{k}(\Pi)$. We further assume bounded variance by restricting $k(\pi,\pi^{\prime})\leq 1$ for all $\pi,\pi^{\prime}\in\Pi$.
This assumption is common in black-box optimization [Srinivas2010] and in repeated games [sessa2020learning, maddux2024multi]. Combined with Assumptions 3 and 4 when $\epsilon^{t}$ is $R$-sub-Gaussian, it allows the leader to learn its unknown cost $J$ using the Gaussian process (GP) framework.
Functions $J\in\mathcal{H}_{k}(\Pi)$ with $\|J\|_{k}<\infty$ can be modeled as a sample from a GP [rasmussen2005, Section 6.2], i.e., $J(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot))$, specified by its mean and covariance functions $\mu(\cdot)$ and $k(\cdot,\cdot)$, respectively. A GP prior $\mathcal{GP}(0,k(\cdot,\cdot))$ over the initial distribution of the unknown function $J$, is used to capture the uncertainty over $J$. Then, for any $\pi\in\Pi$, the function value $J(\pi)$ can be predicated based on a history of measurements $\{y^{\tau}\}_{\tau=1}^{t}$ at points $\{\pi^{\tau}\}_{\tau=1}^{t}$, with $y^{\tau}=J(\pi^{\tau},x^{*}(\pi^{\tau}))+\epsilon^{\tau}$ and $\epsilon^{\tau}\sim\mathcal{N}(0,\sigma^{2})$. Conditioned on the history of measurements, the posterior distribution over $J$ is a GP with mean and variance functions:
$\mu^{t}(\pi)=\textbf{k}^{t}(\pi)^{\top}(\textbf{K}^{t}+\sigma^{2}\textbf{I}^{t})^{-1}\textbf{y}^{t}$ (1e)
$(\sigma^{t})^{2}(\pi)=k(\pi,\pi)-\textbf{k}^{t}(\pi)^{\top}(\textbf{K}^{t}+\sigma^{2}\textbf{I})^{-1}\textbf{k}^{t}(\pi),$ (1f)
where $\textbf{k}^{t}(\pi)=(k(\pi^{\tau},\pi))_{\tau=1}^{t}$, $\textbf{y}^{t}=(y^{\tau})_{\tau=1}^{t}$, and $\textbf{K}^{t}=(k(\pi^{\tau},\pi^{\tau^{\prime}}))_{\tau,\tau^{\prime}=1}^{t}$ is the kernel matrix.
In our setting, however, the leader observes $J(\pi^{t},x^{t}(\pi^{t}))$ rather than $y^{t}$. Thus, we define $\tilde{\mu}^{t}(\cdot)$ which also depends on the true observation as follows:
$\tilde{\mu}^{t}(\pi)=\textbf{k}^{t}(\pi)^{\top}(\textbf{K}^{t}+\sigma^{2}\textbf{I}^{t})^{-1}\mathbf{\tilde{y}}^{t},$ (1g)
where $\mathbf{\tilde{y}}^{t}=(J(\pi^{\tau},x^{\tau}(\pi^{\tau})))_{\tau=1}^{t}$.
IV-A Algorithm and analysis
[…]
$\mathbb{E}[\lVert x^{*}(\pi^{t})-\tilde{x}(\pi^{t})\rVert^{2}]\leq C(\pi^{t})K^{-c}$
for some constants $C(\pi^{t}),c\in\mathds{R}_{+}$. For instance, the works of [drusvyatskiy2022improved, bravo2018bandit] propose such methods that satisfy the above condition under Assumptions 1 and 2.
Upon observing its realized cost $J(\pi^{t},x^{t}(\pi^{t}))$, the leader updates its action by choosing the minimizer of the following surrogate cost function:
$\underline{J}^{t+1}(\pi)\coloneqq\tilde{\mu}^{t}(\pi)-\left(\beta^{t+1}+\frac{\epsilon\sqrt{t+1}}{\sigma}\right)\sigma^{t}(\pi)$
where $\tilde{\mu}^{t}(\cdot)$ and $\sigma^{t}(\cdot)$ are computed as in Equations (1g) and (1f) using past game data $\{(\pi^{\tau},J(\pi^{\tau},x^{\tau}(\pi^{\tau})))\}_{\tau=1}^{t-1}$. Parameter $\beta^{t}$ controls the width of the confidence bound, and if specified adequately, can ensure that $\underline{J}^{t+1}(\pi)$ represents a lower bounds on $J(\pi)$ that holds with high probability.
Algorithm 1
1:Input: $T$, $K$, $\epsilon\geq 0$, $\{\beta^{t}\}_{t=1}^{T}$. Set $\pi^{1}\in\Pi$ randomly.
2:for $t=1,\ldots,T$ do
3:  Leader announces $\pi^{t}$.
4:  Inner loop: Within $K$ rounds, followers compute $\tilde{x}(\pi^{t})=\text{ApproxNE}(\pi^{t},K)$ s.t.
$\mathbb{E}[\lVert x^{*}(\pi^{t})-\tilde{x}(\pi^{t})\rVert^{2}]\leq C(\pi^{t})K^{-c}$ (1h)
5:  Followers set $x^{t}(\pi^{t})=\tilde{x}(\pi^{t})$.
6:  Leader observes $J(\pi^{t},x^{t}(\pi^{t}))$.
7:  Leader updates $\tilde{\mu}^{t+1}$ and $\sigma^{t+1}$ via (1g) and (1f).
8:  Leader chooses:
$\pi^{t+1}\in\displaystyle\arg\min_{\pi\in\Pi}\tilde{\mu}^{t}(\pi)-\left(\beta^{t}+\frac{\epsilon\sqrt{t}}{\sigma}\right)\sigma^{t}(\pi)$
9:end for
We now provide our main result, which shows that the regret of the leader is sublinear in $T$ for properly chosen $\beta^{t}$. Furthermore, if the followers’ utilities are Lipschitz, an approximate Stackelberg equilibrium is learned in time polynomial in $T$ and $K$.
Theorem 1
Let Assumptions 1 - 5 hold, set $\epsilon,\delta\in(0,1)$, and set $\beta^{t}$ equal to $B+(2L_{J}M/\sigma^{2})\sqrt{2(\gamma^{t-1}+1+\log(2/\delta))}$, where the maximum information gain $\gamma^{t-1}$ is a kernel-dependent quantity defined in [Srinivas2010].
  1. 1.
    With probability at least $1-\delta$ the following holds:
    $R^{T}\leq\mathcal{O}((1/\delta+\sqrt{\gamma^{T}})TK^{-\frac{c}{2}}+\beta^{T}\sqrt{\gamma^{T}T})$
    In other words, if $K=\mathcal{O}(T^{\frac{1}{c}})$ and $T=\mathcal{O}\left(\frac{1}{\epsilon^{2}}\right)$ with $c$ set as in Equation (1h) of Algorithm 1, then with probability at least $1-\delta$ it holds that:
    $\min_{t\in[T]}J(\pi^{t},x^{t}(\pi^{t}))$ $\leq J(\pi^{*},x^{*}(\pi^{*}))+\epsilon.$
  2. 2.
    If in addition $U_{i}:\mathcal{X}\times\Pi$ is $L_{U_{i}}$-Lipschitz continuous in $x\in\mathcal{X}$ for all $\pi\in\Pi$ and all $i\in\mathcal{N}$, then with probability at least $1-\delta$ there exists a $t^{*}\in[T]$ such that $(\pi^{t^{*}},x^{t^{*}}(\pi^{t^{*}}))$ is an $\epsilon$-Stackelberg equilibrium.
The regret bound in our theorem depends on two terms: 1) $\mathcal{O}(1/\delta L_{j}K^{-c/2}T)$, which stems from the error accumulation due to non-convergence of the inner-loop to the exact Nash equilibrium, and 2) $\mathcal{O}(\beta^{T}\sqrt{\gamma^{T}T})$, which stems from not knowing the true cost $J$ and estimating it using Gaussian process regression. Note that for some common kernels, explicit bounds on $\gamma^{T}$ are given, e.g., for the linear kernel $\gamma^{T}=\mathcal{O}(d\log T)$ and for the squared exponential kernel $\gamma^{T}=\mathcal{O}(\log^{d+1}T)$ [Srinivas2010, Theorem 5], which depend sublinearly on $T$. By choosing $K$ large enough, i.e., $K=poly(T)$, the approximation error in the equilibrium-finding subroutine becomes negligible and our theorem shows that with high probability the leader finds an action that achieves nearly optimal cost function value in a sublinear number of rounds $T$. We now provide a proof of our theorem.
Proof:
We start by proving part 1) of the theorem. The regret of the leader is upper-bounded as follows:
$R^{T}$ $=\sum_{t=1}^{T}J(\pi^{t},x^{t}(\pi^{t}))-\min_{\pi\in\Pi}J(\pi,x^{*}(\pi))$
$\leq\sum_{t=1}^{T}\lvert J(\pi^{t},x^{t}(\pi^{t}))-J(\pi^{t},x^{*}(\pi^{t}))\rvert$
$+\sum_{t=1}^{T}J(\pi^{t},x^{*}(\pi^{t}))-\min_{\pi\in\Pi}J(\pi,x^{*}(\pi)).$ (1i)
Then, due to Assumption 4, we have:
$R^{T}$ $\leq\underbrace{\sum_{t=1}^{T}L_{J}\lVert x^{*}(\pi^{t})-x^{t}(\pi^{t})\rVert}_{\coloneqq\Delta_{F}}$
$+\underbrace{\sum_{t=1}^{T}J(\pi^{t},x^{*}(\pi^{t}))-\min_{\pi\in\Pi}J(\pi,x^{*}(\pi))}_{\coloneqq\Delta_{L}}.$ (1j)
We start by bounding the term $\Delta_{F}$. Note that by Jensen’s inequality and Line (1h) in Algorithm 1 the following holds:
$\bigl(\mathbb{E}\bigl[\lVert x^{*}(\pi^{t})-x^{t}(\pi^{t})\rVert\bigr]\bigr)^{2}$ $\leq\mathbb{E}\bigl[\lVert x^{*}(\pi^{t})-x^{t}(\pi^{t})\rVert^{2}\bigr]$
$\leq C(\pi^{t})K^{-c}$ $\leq CK^{-c},$ (1k)
where in the last inequality we used that $\Pi$ is compact, which ensures the existence of a maximizer $C=\max_{\pi\in\Pi}C(\pi)$. Furthermore, first applying Markov’s inequality and then Inequality (1k) yields:
$\mathbb{P}(\Delta_{F}$ $\geq\frac{2}{\delta}TL_{J}\sqrt{C}K^{-\frac{c}{2}})$
$\leq\frac{\delta\sum_{t=1}^{T}\mathbb{E}\bigl[\lVert x^{*}(\pi^{t})-x^{t}(\pi^{t})\rVert\bigr]}{2T\sqrt{C}K^{-\frac{c}{2}}}\leq\frac{\delta}{2}.$ (1l)
Next, we bound the term $\Delta_{L}$, which corresponds to the regret $R^{T}$ of the leader assuming the lower-level game were at an exact Nash equilibrium $x^{*}(\pi^{t})$. [^2] Recall that at the exact Nash equilibrium $x^{*}(\pi^{t})$ the leader’s cost is uniquely defined by $J(\pi^{t}):=J(\pi^{t},x^{*}(\pi^{t}))$. Note that by Assumption 5 the cost function $J$ has a bounded RKHS norm. If the leader further observed $y^{t}=J(\pi^{t},x^{*}(\pi^{t}))+\epsilon^{t}$, where $\epsilon^{t}$ is zero-mean and $R$-sub-Gaussian noise, then $\Delta_{L}$ could be upper-bounded leveraging the Gaussian process framework, for example by [chowdhury2017kernelized, Theorem 3]. In our setting, however, the zero-mean assumption on the noise is not necessarily satisfied. Namely, the leader observes $J(\pi^{t},x^{t}(\pi^{t}))$ at each round $t$ rather than $J(\pi^{t},x^{*}(\pi^{t}))$. The error term $\epsilon^{t}=J(\pi^{t},x^{*}(\pi^{t}))-J(\pi^{t},x^{t}(\pi^{t}))$, resulting from approximating the Nash equilibrium, is not guaranteed to be zero-mean.
To alleviate this, we instead rewrite $\epsilon^{t}$ as: ϵt = E[J(πt, x*(πt)) - J(πt,xt(πt))] + ~ϵt, where the expectation is taken with respect to the randomness in the inner-loop for finding $x^{t}(\pi^{t})$, $\tilde{\varepsilon}^{t}$ is zero mean by construction, and $2L_{J}M$-sub-Gaussian as Assumption 4 ensures the distribution of $\tilde{\epsilon}^{t}$ is bounded in $2L_{J}M,2L_{J}M]$.
We furthermore define a new cost function $\tilde{J}:\Pi\rightarrow\mathds{R}$ as $\tilde{J}(\pi,x^{*}(\pi))=J(\pi,x^{*}(\pi))+\mathbb{E}[J(\pi,\tilde{x}(\pi))-J(\pi,x^{*}(\pi))]$, where $\tilde{x}(\pi)=\text{ApproxNE}(\pi,K)$ in Algorithm 1, which corresponds to the true cost function at the Nash equilibrium plus the expected error from being at an approximation of the Nash equilibrium. Now the leader’s observation $J(\pi^{t},x^{t}(\pi^{t}))$ satisfies $J(\pi^{t},x^{t}(\pi^{t}))=\tilde{J}(\pi^{t},x^{*}(\pi^{t}))+\tilde{\epsilon}^{t}$, where $\tilde{\epsilon}^{t}$ is zero-mean, $2L_{J}M$-sub-Gaussian, but $\tilde{J}$ may no longer satisfy Assumption 5, namely $\tilde{J}\not\in\{f\in\mathcal{H}_{k}(\Pi)\mid\lVert f\rVert_{k}\leq B\}$. In the literature, this setting is referred to as the misspecified setting [bogunovic2021misspecified]. Note however that $\tilde{J}$ can be uniformly approximated by $J$ as follows:
$\lVert\tilde{J}-J\rVert_{\infty}$ $=\sup_{\pi\in\Pi}\lvert\tilde{J}(\pi,x^{*}(\pi))-J(\pi,x^{*}(\pi))\rvert$
$=\sup_{\pi\in\Pi}\lvert\mathbb{E}[J(\pi,\tilde{x}(\pi))-J(\pi,x^{*}(\pi))]\rvert$
$\overset{(i)}{\leq}L_{J}\sup_{\pi\in\Pi}\mathbb{E}[\lVert x^{*}(\pi)-\tilde{x}(\pi)\rVert]$
$\overset{(ii)}{\leq}L_{J}\sup_{\pi\in\Pi}\sqrt{\mathbb{E}[\lVert x^{*}(\pi)-\tilde{x}(\pi)\rVert^{2}]}$
$\overset{(iii)}{\leq}L_{J}C^{\frac{1}{2}}K^{-\frac{c}{2}},$
where in $(i)$ we used Assumption 4, in $(ii)$ we applied Jensen’s inequality, and in $(iii)$ we used Equation (1h). In summary, the assumptions of [bogunovic2021misspecified, Theorem 1] are now satisified as $\tilde{\epsilon}^{t}$ is sub-Gaussian, $J\in\{f\in\mathcal{H}_{k}(\Pi)\mid\lVert f\rVert_{k}\leq B\}$, and $\tilde{J}$ can be uniformly approximtaed by $J$. Then, since the leader chooses $\pi^{t}$ via Equation (8), by [bogunovic2021misspecified, Theorem 1], with probability at least $1-\delta/2$ it holds that:
$\Delta_{L}=2\sqrt{(2\sigma^{2}+1)\gamma^{T}}\big(\beta^{T}\sqrt{T}+T\frac{L_{J}C^{\frac{1}{2}}K^{-\frac{c}{2}}}{\sigma}\big)$ (1m)
Next we bound $R^{T}$ leveraging the bounds we established for $\Delta_{F}$ and $\Delta_{L}$. To this end, let $E_{F}$ and $E_{L}$ denote the events:
$E_{F}$ $=\big(\Delta_{F}\geq\frac{2}{\delta}L_{J}TC^{\frac{1}{2}}K^{-\frac{c}{2}}\big)$
$E_{L}$ $=\big(\Delta_{L}\geq 2\sqrt{(2\sigma^{2}+1)\gamma^{T}}\big(\beta^{T}\sqrt{T}+T\frac{L_{J}C^{\frac{1}{2}}K^{-\frac{c}{2}}}{\sigma}\big)$
and denote by $\overline{E}_{F}$ and $\overline{E}_{L}$ their corresponding complements. Then, it follows that:
$\mathbb{P}\big(R^{T}\leq\left(\frac{2}{\delta}+\frac{2\sqrt{(2\sigma^{2}+1)\gamma^{T}}}{\sigma}\right)L_{J}TC^{\frac{1}{2}}K^{-\frac{c}{2}}+$
$\quad\quad 2\beta^{T}\sqrt{\gamma^{T}T(2\sigma^{2}+1)}\big)$
$\geq\mathbb{P}\big(\big(\Delta_{F}\leq\frac{2}{\delta}L_{J}TC^{\frac{1}{2}}K^{-\frac{c}{2}}\big)$
$\quad\quad\cap\big(\Delta_{L}\leq 2\sqrt{(2\sigma^{2}+1)\gamma^{T}}(\beta^{T}\sqrt{T}+T\frac{L_{J}C^{\frac{1}{2}}K^{-\frac{c}{2}}}{\sigma}\big)\big)$
$=\mathbb{P}\left(\overline{E}_{F}\cap\overline{E}_{L}\right)$
$=1-\mathbb{P}\bigl(E_{F}\cup E_{L}\bigr)$
$\geq 1-\bigl(\mathbb{P}(E_{F})+\mathbb{P}(E_{L})\bigr)$
$=1-(\frac{\delta}{2}+\frac{\delta}{2}).$
where we used the union bound in the second-to-last line and (1l) and (1m) in the last line. Finally, we have $\mathbb{P}\big(R^{T}\leq\mathcal{O}((\tfrac{1}{\delta}+\sqrt{\gamma^{T}})TK^{-\tfrac{c}{2}}+\beta^{T}\sqrt{T\gamma^{T}})\big)\geq 1-\delta$.
Now, we proceed to prove part 2) of the theorem. Let $\pi^{*}\in\arg\min_{\pi\in\Pi}J(\pi,x^{*}(\pi))$. It can be easily verified that $(\pi^{*},x^{*}(\pi^{*}))$ is a Stackelberg equilibrium. Set $t^{*}\in\arg\min_{t\in[T]}J(\pi^{t},x^{t}(\pi^{t}))$, then, by definition of $R^{T}$ it follows that with probability at least $1-\delta$:
$J(\pi^{t^{*}},x^{t^{*}}(\pi^{t^{*}}))$ $\leq J(\pi^{*},x^{*}(\pi^{*}))+2\frac{\beta^{T}}{\sqrt{T}}\sqrt{\gamma^{T}(2\sigma^{2}+1)}\big)$ (1n)
$+\left(\frac{2}{\delta}+\frac{2\sqrt{(2\sigma^{2}+1)\gamma^{T}}}{\sigma}\right)L_{J}C^{\frac{1}{2}}K^{-\frac{c}{2}}$ (1o)
Set $K$ and $T$ as follows:
$K=\mathcal{O}(T^{\frac{1}{c}})$ (1p)
$T=\mathcal{O}\bigg(\frac{1}{\epsilon^{2}}\max\big\{\displaystyle\max\{1,L_{U}^{2}\}\frac{L_{J}^{2}C}{\delta^{2}},\max\{(\beta^{T})^{2},\frac{L_{j}^{2}C}{\sigma^{2}}\}\gamma^{T}\big\}\bigg)$ (1q)
where $L_{U}=\max_{i\in\mathcal{N}}L_{U_{i}}$. Then, by plugging the values of $K$ and $T$ into Equation (1n) it follows that:
$J(\pi^{t^{*}},x^{t^{*}}(\pi^{t^{*}}))$ $\leq J(\pi^{*},x^{*}(\pi^{*}))+\epsilon$
$\leq J(\overline{\pi},x^{*}(\overline{\pi}))+\epsilon,\quad\forall\overline{\pi}\in\Pi,$
where $x^{*}(\overline{\pi})$ is the unique Nash equilibrium of the game $\Gamma(\mathcal{N},\{\mathcal{X}_{i}\}_{i=1}^{N},\{U_{i}\}_{i=1}^{N};\overline{\pi})$. Furthermore, if in addition $U_{i}:\mathcal{X}\times\Pi\rightarrow\mathds{R}$ is $L_{U_{i}}$-Lipschitz continuous in $x\in\mathcal{X}$ for all $\pi\in\Pi$ and all $i\in\mathcal{N}$, then the following holds:
$\mathbb{P}\bigg(\lvert U_{i}(x^{*}(\pi^{t^{*}});\pi^{t^{*}})-U_{i}(x^{t^{*}}(\pi^{t^{*}});\pi^{t^{*}})\lvert\geq\frac{\displaystyle L_{U_{i}}\sqrt{C}K^{-\frac{c}{2}}}{\delta}\bigg)$
$\leq\mathbb{P}\bigg(\lVert x^{*}(\pi^{t^{*}})-x^{t^{*}}(\pi^{t^{*}})\lVert\geq\frac{\sqrt{C}K^{-\frac{c}{2}}}{\delta}\bigg)$
$\leq\frac{\delta\mathbb{E}\big[\lVert x^{*}(\pi^{t^{*}})-x^{t^{*}}(\pi^{t^{*}})\rVert\big]}{\sqrt{C}K^{-\frac{c}{2}}}\leq\delta,$
where in the last inequality we used Markov’s inequality and Line (1h) in Algorithm 1. Thus, with $K$ and $T$ set as in Equations (1p) and (1q), respectively, with probability at least $1-\delta$ it holds that:
$U_{i}(x^{t^{*}}(\pi^{t^{*}});\pi^{t^{*}})+\epsilon$ $\geq U_{i}(x^{*}(\pi^{t^{*}});\pi^{t^{*}})$
$\geq U_{i}(x_{i},x_{-i}^{*}(\pi^{t^{*}});\pi^{t^{*}}),\quad\forall x_{i}\in\mathcal{X}_{i},$
where we used that $x^{*}(\pi^{t^{*}})$ is the unique Nash equilibrium of $\Gamma(\mathcal{N},\{\mathcal{X}_{i}\}_{i=1}^{N},\{U_{i}\}_{i=1}^{N};\pi^{t^{*}})$. We conclude that $(\pi^{t^{*}},x^{t^{*}}(\pi^{t^{*}}))$ is an $\epsilon$ -Stackelberg equilibrium.
For a single-leader multi-follower Stackelberg game, [maheshwari2023follower] shows that with $K=\mathcal{O}(T^{\frac{1}{2c}})$ and $T=\mathcal{O}(1/\epsilon^{2})$, their algorithm converges to an $\epsilon$-stationary point, i.e., $\min_{t\in[T]}\mathbb{E}[\lVert\nabla J(\pi^{t})\rVert^{2}]\leq\epsilon$. While their convergence time matches ours up to constant terms, their approach requires a hard-to-verify assumption on the Hessian of the leader’s cost function to additionally ensure convergence to an approximate Stackelberg equilibrium. Moreover, their method requires two-point feedback to estimate the hypergradient of the leader’s objective while our algorithm relies on one-point feedback to estimate the leader’s cost function.
[…]
Figure 1: Illustration of the setup with 2 districts and 3 ride-hailing companies.
V Coordination of ride-hailing companies: Numerical study
We apply Algorithm 1 to the pricing problem in electric ride-hailing markets, introduced in Section III. In our numerical study illustrated in Figure 1, for the lower level, we consider a city region divided into two districts, the outskirts, and a downtown area, where $|\mathcal{N}|=3$ competing ride-hailing companies operate and can recharge their fleets of scaled sizes $M=(2,4,6)$. To reflect the typical higher demand in downtown areas, we set the total revenue potential and abandonment vectors in Equation (1b) to $W=(30,60)$ and $\Delta=(0.1,0.5)$, respectively.
In this simplified setup, we assume that the regulatory authority aims to select an optimal charging price vector $\pi\in\mathds{R}^{2}$ within the range $[0.1,5.0]$ to help balance the demand on the power grid and reduce the number of idling vehicles in the downtown area. Specifically, to counterbalance the increased attractiveness of the downtown district, the authority aims to set lower charging prices in the outskirts, with the goal of steering the vehicle distribution toward a uniform spread, i.e., $\xi^{*}=(0.5,0.5)$. For example, if the central authority sets prices proportional to the revenue potential of the districts, e.g., $\pi_{\text{base}}=(1,2)$, then the attained leader’s cost equals $J(\pi_{\text{base}})\approx 0.22$ while $J(\pi^{*})=0$ if the central authority sets prices optimally.
We implement Algorithm 1 by choosing a squared exponential kernel, combined with a standard heuristic approach from [brochu2009tutorial] to approximate the minimum of the optimistic estimate of the cost function in Line 8 of Algorithm 1. To demonstrate the practicality of our learning method, we evaluate its performance under various levels of approximation error when solving the lower-level Nash equilibrium. Specifically, we terminate the inner loop when $\left\lVert x^{*}(\pi^{t})-x^{t}(\pi^{t})\right\rVert\leq\varepsilon$ for $\varepsilon\in\{10^{-6},0.1,0.3,0.5\}$. For each $\varepsilon$, the GP parameters are calibrated via standard gradient-based optimization of the marginal log-likelihood [rasmussen2005], using data collected after $N_{\text{warm}}=5$ random iterations of the outer loop. While the value of $\beta^{T}$ proposed by Theorem 1 provides theoretical guarantees, in practice we empirically found that fixing $\beta^{T}=0.2$ enables the leader to find a high-quality pricing for all $\varepsilon$ values, while avoiding the computational burden of finding the problem-specific constants required by Theorem 1.
The results are illustrated in Figure 2 and further supported by numerical values in Table I.
[…]
Figure 2: The top figure illustrates the average cumulative regret of the regulatory authority, $R^{t}/t$, while the bottom figure displays the leader’s objective, both over $T=25$ iterations. The initial $N_{\text{warm}}=5$ iterations correspond to a warm-up phase, during which pricing vectors are selected randomly in order to collect data for calibrating hyperparameters of the GP. Different colors represent varying levels of approximation error in computing the Nash equilibrium within the inner loop of Algorithm 1.
TABLE I: Charging prices and leader’s performance
Approximation
error $\varepsilon$
Best pricing vector $J(\pi^{t},x^{t}(\pi^{t}))$ $R^{t}/t$
$\pi_{1}^{t}$ $\pi_{2}^{t}$
$10^{-6}$ 1.9157 4.99 $6.3\cdot 10^{-5}$ 0.0595
$0.1$ 1.9134 4.99 $6.2\cdot 10^{-5}$ 0.0590
$0.3$ 1.8761 4.99 $9.2\cdot 10^{-4}$ 0.0572
$0.5$ 1.8305 4.99 $1.3\cdot 10^{-3}$ 0.0679
The cumulative regret plot suggests that our proposed framework is fairly robust to approximation errors in the inner loop, as all curves for $\varepsilon\leq 0.3$ show a similar downward trend over time, with the $\varepsilon=10^{-6}$ and $\varepsilon=0.1$ curves almost perfectly overlapping. For $\varepsilon=0.5$, the trend remains consistent but decreases at a slightly slower rate. Interestingly, the heuristic from [brochu2009tutorial], when combined with a fixed $\beta^{T}$, led to the fastest reduction in average cumulative regret and the most rapid initial learning for $\varepsilon=0.3$. However, the lowest leader’s objective is obtained when the approximation error is negligible.
VI Conclusion
In this paper, we studied the problem of learning to play single-leader, multi-follower Stackelberg games when the lower-level game is unknown to the leader. We proposed a novel no-regret algorithm, and proved that it converges to an $\epsilon$-Stackelberg equilibrium in $\mathcal{O}(\sqrt{T})$ rounds with high probability under kernel-based regularity assumptions. Our method improves practicality by removing the need for gradient-based techniques that require access to followers’ private utilities. Lastly, we validated our method in a numerical case study on electricity pricing, demonstrating its convergence under varying levels of lower-level approximation error. Future work may explore extensions to contextual settings, where the lower-level game depends on both the leader’s action and a random context, accounting for external factors like weather and time that are beyond the leader’s control.